# Photometric classification of stars around the Milky Way’s central black hole

### v1

@@ 7.5 IMF of early-type stars
To infer the IMF of the young cluster formed with the early-type candidates brighter than 16 magnitudes and located within a distance smaller than the computed break radius of 9′′, we employed an optimization algorithm that compares observational data with theoretical simulations. Utilizing SPISEA (Hosek Jr et al., 2020), we constructed a grid of synthetic single-age star clusters varying in age, cluster mass, extinction, and IMF slope index. To generate synthetic photometry for each star, we utilized the NACO KS filter, assuming a distance of 8.25 kpc and a metallicity of [M/H] = 0.15. SPISEA allows us to consider multiplicity in the generated clusters by constructing a multiplicity object (multiplicity.MultiplicityUnresolved) based on Lu et al. (2013). This configuration treats star systems as unresolved, combining all components into a single star within the cluster. Due to the absence of stars with lower masses in our observed magnitude ranges, the mass limits for stars in the simulated clusters were set to 5 Mdirect-$\odot$-120 Mdirect-$\odot$. We used the MIST (MESA Isochrones & Stellar Tracks; Choi et al., 2016) evolution model and adopted the reddening law from Nishiyama et al. (2009). We examined isochrones spanning a range of cluster ages from 2 to 8 Myr and constructed synthetic KLFs for each combination of parameters. Subsequently, we compared the theoretical KLFs to the observed KLF by computing the chi-squared statistic. Finally, we identified the parameter set that minimizes the chi-squared value. The optimal fit corresponds to a cluster age of 4 Myr. We utilized a Monte Carlo (MC) approach to incorporate the uncertainty arising from the sampling of the IMF in each realization of a cluster with SPISEA. It is important to note that the choice of the evolution model can impact the fit values, particularly the inferred age. The best power-law IMF index value is $1.6\pm 0.1$, obtained for the three sets of early-type candidates using different methods. To validate our results, we also fitted the IMF by limiting the magnitude range up to 15. In this analysis, we determined that the IMF slope is approximately 1.7, falling within the expected range and accounting for uncertainties in our results.
@@
Figure 20: Comparison between different models simulated with SPISEA and the observed KLF for the different methods summing up probabilities and corrected by completeness (see Section 6.2)
. The red line represents the KLF corresponding to a simulated cluster with the optimal power-law IMF, featuring an IMF slope of -1.6, while the shaded region indicates the uncertainty in the model. Additionally, the KLFs of cluster models computed with SPISEA using the Kroupa IMF (Kroupa, 2001) and employing a top-heavy IMF (Bartko et al., 2010) are presented for comparison.
@@ 8 Conclusions
We have classified stars within the central parsec around Sgr A*, utilizing NIR IB images obtained with NACO/VLT. While these data have been previously investigated, we revisit their analysis, improving the reduction and photometric processes, leading to enhanced sensitivity and more reliable estimates of photometric uncertainties compared to prior efforts. Additionally, we employ three distinct yet complementary classification methods for distinguishing stars as early or late type: Bayesian inference, a basic neural network, and a fast gradient-boosted trees algorithm. Thanks to the information from extensive spectroscopic studies conducted over the past decade, we expand our database, enhancing the training dataset for our models and thereby obtaining more robust and confident results.
Despite challenges associated with studying this region, such as extreme interstellar extinction and crowding, we identified 155 new early star candidates using Bayesian and MLP methods. The identified candidates present promising targets for upcoming spectroscopic observations, which are crucial for validating their classification and providing additional support for our results. These candidates, while intriguing as potential spectroscopic targets, are not the primary outcome of our probabilistic approaches.
Consistent with previous work we have found that broken power laws provide adequate fits to the radial density distribution of the late and early-type stars. Our investigation reveals a break radius of $7.4\pm 1.2^{\prime\prime}$ for the late-type distribution and $9.2\pm 0.6^{\prime\prime}$ for the early-type distribution. The slopes for the late-type distribution are determined to be $\beta=0.46\pm 0.07$ (outside of the break radius) and $\gamma=0\pm 0.08$, while the early-type distribution exhibits slopes of $\beta=3.65\pm 0.47$ and $\gamma=0.81\pm 0.08$.
@@
We found that the mass function of the early type stars appears to be top-heavy near Sgr A*, with a power-law slope of $1.6\pm 0.1$. Contrary to Bartko et al. (2010) we found no peak of the luminosity function around 14 mag and a significantly steeper mass function. This is probably related to the incompleteness of the early spectroscopic data at faint magnitudes.
At larger projected distances, the best-fit models varied significantly among the methods, with all IMF slopes exceeding 1.6 and possibly in agreement with a standard Salpeter/Kroupa IMF. These findings suggest distinct star formation mechanisms between the inner and outer regions, with deviations from the standard Salpeter/Kroupa IMF indicating unique processes in the central region and a potentially evolving environment further out.
@@ Appendix E Exploring specific stars
In this section, we examine several stars of particular interest.
Firstly, we analyzed five stars with conflicting spectral classifications and no apparent basis for preferring one over the other. We choose their types based on their positions in the CBDD (see Fig. 29). The misclassification could be attributed to incorrect catalogue matching. We favour the late classification for only one star, our Id 1576, which corresponds to Id 553 in Feldmeier-Krause et al. (2017a), in contrast to the early classification assigned by Yelda et al. (2014) as S13-3, based on the reuse of data from Paumard et al. (2006). We note that this star is not categorized as early in von Fellenberg et al. (2022), the most recent MPE group work. The probabilities support this classification with a confidence level of 1.4% or less for being classified as early.
@@
As we approached the conclusion of our study, we became aware of the work by Jia et al. (2023). After review, we found that four out of the six “new” early-type stars we identified correspond to stars in the von Fellenberg et al. (2022) catalogue and, to some extent, in the Yelda et al. (2014) catalogue. Specifically, the stars S11-246 and S5-106 are spectroscopic unclassified in our combined catalogue (designated as Our Id 3588 and 5142). Our early probabilities range from 76% to 94%, depending on the star and method used, thus affirming the validity of our classification methods.
@@ Appendix G Bayesian fits
In this section, we include other use figures for the Bayesian analysis.
Firstly that is the likelihood, prior and posterior in magnitude radial space in Fig. 30. We show the KLF of the likelihood in Figure 31.
@@ Appendix H Multi-layer Perceptron results
More important results are shown in Fig. 35.
@@ Appendix I Gradient boosted trees results
More important results are shown in Fig. 36.
@@
Figure 37: SEDs for early-type candidates detected only with MLP method, located in the upper left corner (left panel), lower left corner (middle panel), and upper right corner (right panel).

### v2

@@ 7.5 IMF of early-type stars
To infer the IMF of the young cluster formed with the early-type candidates brighter than 16 magnitudes and located within a distance smaller than the computed break radius of 9′′, we employed an optimization algorithm that compares observational data with theoretical simulations. Utilizing SPISEA (Hosek Jr et al., 2020), we constructed a grid of synthetic single-age star clusters varying in age, cluster mass, extinction, and IMF slope index. To generate synthetic photometry for each star, we utilized the NACO KS filter, assuming a distance of 8.25 kpc and a metallicity of [M/H] = 0.15. SPISEA allows us to consider multiplicity in the generated clusters by constructing a multiplicity object (‘multiplicity.MultiplicityUnresolved’) based on Lu et al. (2013). This configuration treats star systems as unresolved, combining all components into a single star within the cluster. Due to the absence of stars with lower masses in our observed magnitude ranges, the mass limits for stars in the simulated clusters were set to 5 Mdirect-$\odot$-120 Mdirect-$\odot$. We used the MIST (MESA Isochrones & Stellar Tracks; Choi et al., 2016) evolution model and adopted the reddening law from Nishiyama et al. (2009). We examined isochrones spanning a range of cluster ages from 2 to 8 Myr and constructed synthetic KLFs for each combination of parameters. Subsequently, we compared the theoretical KLFs to the observed KLF by computing the chi-squared statistic. Finally, we identified the parameter set that minimizes the chi-squared value. The optimal fit corresponds to a cluster age of 4 Myr. We utilized a Monte Carlo (MC) approach to incorporate the uncertainty arising from the sampling of the IMF in each realization of a cluster with SPISEA. It is important to note that the choice of the evolution model can impact the fit values, particularly the inferred age. The best power-law IMF index value is $1.6\pm 0.1$, obtained for the three sets of early-type candidates using different methods. To validate our results, we also fitted the IMF by limiting the magnitude range up to 15. In this analysis, we determined that the IMF slope is approximately 1.7, falling within the expected range and accounting for uncertainties in our results.
@@
Figure 20: Comparison between different models simulated with SPISEA and the observed KLF for the different methods summing up probabilities and corrected by completeness (see Section 6.2). The red line represents the KLF corresponding to a simulated cluster with the optimal power-law IMF, featuring an IMF slope of -1.6, while the shaded region indicates the uncertainty in the model. Additionally, the KLFs of cluster models computed with SPISEA using the Kroupa IMF (Kroupa, 2001) and employing a top-heavy IMF (Bartko et al., 2010) are presented for comparison.
@@ 8 Conclusions
In this paper, we classify stars within the central parsec around Sgr A*, utilizing NIR IB images obtained with NACO/VLT. While these data have been previously investigated, we revisit their analysis, improving the reduction and photometric processes, leading to enhanced sensitivity and more reliable estimates of photometric uncertainties compared to prior efforts. Additionally, we employ three distinct yet complementary classification methods for distinguishing stars as early or late type: Bayesian inference, a basic neural network, and a fast gradient-boosted trees algorithm. Thanks to the information from extensive spectroscopic studies conducted over the past decade, we expand our database, enhancing the training dataset for our models and thereby obtaining more robust and confident results.
Despite challenges associated with studying this region, such as extreme interstellar extinction and crowding, we identify 155 new early star candidates using Bayesian and MLP methods. The identified candidates present promising targets for upcoming spectroscopic observations, which are crucial for validating their classification and providing additional support for our results. These candidates, while intriguing as potential spectroscopic targets, are not the primary outcome of our probabilistic approaches.
Consistent with previous work we find that broken power laws provide adequate fits to the radial density distribution of the late and early-type stars. Our investigation reveals a break radius of $7.4\pm 1.2^{\prime\prime}$ for the late-type distribution and $9.2\pm 0.6^{\prime\prime}$ for the early-type distribution. The slopes for the late-type distribution are determined to be $\beta=0.46\pm 0.07$ (outside of the break radius) and $\gamma=0\pm 0.08$, while the early-type distribution exhibits slopes of $\beta=3.65\pm 0.47$ and $\gamma=0.81\pm 0.08$.
@@
We find that the mass function of the early type stars appears to be top-heavy near Sgr A*, with a power-law slope of $1.6\pm 0.1$. Contrary to Bartko et al. (2010) we find no peak of the luminosity function around 14 mag and a significantly steeper mass function. This is probably related to the incompleteness of the early spectroscopic data at faint magnitudes.
At larger projected distances, the best-fit models vary significantly among the methods, with all IMF slopes exceeding 1.6 and possibly in agreement with a standard Salpeter/Kroupa IMF. These findings suggest distinct star formation mechanisms between the inner and outer regions, with deviations from the standard Salpeter/Kroupa IMF indicating unique processes in the central region and a potentially evolving environment further out.
@@ Appendix E Exploring specific stars
In this section, we examine several stars of particular interest. Firstly, we analyzed five stars with conflicting spectral classifications and no apparent basis for preferring one over the other. We choose their types based on their positions in the CBDD (see Fig. 29). The misclassification could be attributed to incorrect catalogue matching. We favour the late classification for only one star, our Id 1576, which corresponds to Id 553 in Feldmeier-Krause et al. (2017a), in contrast to the early classification assigned by Yelda et al. (2014) as S13-3, based on the reuse of data from Paumard et al. (2006). We note that this star is not categorized as early in von Fellenberg et al. (2022), the most recent MPE group work. The probabilities support this classification with a confidence level of 1.4% or less for being classified as early.
@@
As we approached the conclusion of our study, we became aware of the work by Jia et al. (2023). After review, we found that four out of the six ‘new’ early-type stars we identified correspond to stars in the von Fellenberg et al. (2022) catalogue and, to some extent, in the Yelda et al. (2014) catalogue. Specifically, the stars S11-246 and S5-106 are spectroscopic unclassified in our combined catalogue (designated as Our Id 3588 and 5142). Our early probabilities range from 76% to 94%, depending on the star and method used, thus affirming the validity of our classification methods.
@@ Appendix G Bayesian fits
In this section, we present additional figures for the Bayesian analysis. First, we show the likelihood, prior, and posterior in magnitude-radial space in Fig. 30. We show the KLF of the likelihood in Figure 31.
@@ Appendix H Multi-layer perceptron results
Figure  35 shows the probabilities of the stars obtained by the MLP method in CBD-magnitude-radius space, where the radius is the distance from Sgr A*.
@@ Appendix I Gradient boosted trees results
Figure  36 shows the probabilities obtained by the random forest in CBD-magnitude-radius space.
@@
Figure 37: SEDs for early-type candidates detected only with MLP method, located in the upper left corner (upper panel), lower left corner (middle panel), and upper right corner (bottom panel).